OpenStudy (anonymous):

can someone please solve and explain the following quadratic equation? 4x^3 - 3x^3 + 60x = 0, x=?

OpenStudy (anonymous):

This is not a quadratic equation!!!

OpenStudy (anonymous):

Its cubic

OpenStudy (anonymous):

check if u hv typed it properly

OpenStudy (anonymous):

I did type it properly but it says in my book it is a quadratic equation;/

OpenStudy (anonymous):

book is lying

OpenStudy (anonymous):

@Harkirat what was with all the exclamation points?

OpenStudy (anonymous):

There can be typographical errors in a book.....

OpenStudy (anonymous):

Well, assuming that there was a typo, and the equation is as follow \[x ^{3} - 3x^{2} + 60x = 0 \] we can do the following: \[x(x^{2} - 3x + 60) = 0\] \[x=0\] is one answer , the other ones, just solve \[x^{2} - 3x + 60 = 0\]

OpenStudy (anonymous):

@Dijah Just expressing my surprise .. :))

OpenStudy (anonymous):

well if you simplify it, it becomes x(x^2 +60) = 0, and then you would simply be solving a qudratic

OpenStudy (anonymous):

@poring she has not confirmed the typo....

OpenStudy (anonymous):

is there anyway to solve this problem though? I couldn't figure it out...

OpenStudy (anonymous):

@Lydia Is the second term -3x^3 or -3x^2 ??

OpenStudy (anonymous):

4x^3 - 3x^3 +60x = 0 simplifies to x^3 +60x = 0. Then factor out an x and you have x(x^2+60) = 0. Then one solution of x will be 0, and you will get the other two solutions by applying the quadratic formula to x^2 +60 = 0.

OpenStudy (anonymous):

@Hakirat it's -3x^3

OpenStudy (anonymous):

if it helps i have the answer, i just don't understand how they got it....

OpenStudy (radar):

Assuming there is no typo:\[4x ^{3}-3x ^{3}+60x=0\] \[x ^{3}+60x=0\] \[x(x ^{2}+60)=0\] x=0 \[x ^{2}+60=0\] \[x=\sqrt{-60}=\sqrt{4}\sqrt{-15}=2\sqrt{15}i\]

OpenStudy (anonymous):

thank you @radar, but the book must have had a typo...because the answer in the book says (0,5,3), and there's no possible way to get that unless the book actually typed the equation wrong.

OpenStudy (radar):

I really think they meant for the second term to be some x^2 rather than the -3x^3.

OpenStudy (anonymous):

i think you're right, because all the other equations in the book had the second x raised to the 2nd power except this one...plus the answer doesn't match up. but thanks for your guys' help:)

OpenStudy (anikate):

what is the question?

OpenStudy (anikate):

if its an equation that i no how to do. I will help

OpenStudy (radar):

you are welcome, one other thing, I don't know how they have been teaching, but the square root of 4 is actually +/1 2 so the answer might also be \[\pm 2\sqrt{15}i\]

OpenStudy (anikate):

no its a simple 2

OpenStudy (anonymous):

thank you radar!! :D

OpenStudy (radar):

\[*** \pm 2\]

OpenStudy (anonymous):

kool

OpenStudy (radar):

It would have three roots: 0, and\[2\sqrt{15}i, and -2\sqrt{15}i\]

OpenStudy (anonymous):

@radar is right

OpenStudy (anonymous):

I worked backwards from the three given roots and landed up with the equation x^3 - 8x^2 + 15x = 0 multiplying throughout by 4, we get 4x^3 - 32x^2 + 60x = 0 Anybody for confirming this ????

OpenStudy (radar):

Was this the roots of 0, 5, and 3?

OpenStudy (anonymous):

yes, that is what she mentioned......I think.......

OpenStudy (radar):

x=0 for sure lol, I think that was what was wanted to be solved, but it got typoed!

OpenStudy (radar):

@harikat, the 5 works also

OpenStudy (radar):

Confirmed

OpenStudy (anonymous):

so we find the right equation for these values is 4x^3 - 32x^2 + 60x = 0 So Lydia, I suggest you modify the equation to what i hv written and solve it you will get the roots given as answer.......

OpenStudy (anonymous):

@Harkirat your picture reminds me of the lion king

OpenStudy (anonymous):

Thanks for liking it. Actually I am a Sikh from the Punjab state of India. My community is lion-hearted by nature and my surname also means "lion". So I found this nice picture and used it....

OpenStudy (anonymous):

I am thankful to all who gave me medals o((:-})>

OpenStudy (anonymous):

it was me ;)

OpenStudy (anonymous):

thank u very much... ☺

OpenStudy (anonymous):

thank you guys for helping me with this, it really saved me some worrying:) good teamwork!